A plane can fly 340mph in still air. If it can fly 200 miles downwind in the same amount of time it can fly 140 miles upwind, find the velocity of the wind.
Answer by mananth(16949)

A plane can fly 340mph in still air. If it can fly 200 miles downwind in the same amount of time it can fly 140 miles upwind, find the velocity of the wind.
..
let speed of wind be x mph
..
speed downwind = 340+x
speed upwind = 340-x
..
time downwind = 200/340+x
time upwind = 140/340-x
..
200/(340+x)= 140/(340-x)
200(340-x)=140(340+x)
68000-200x= 47600+140x
68000-47600 = 200+140x
20400 = 340x
x= 20400/34
x= 60mph the speed of wind